A bacterial population is known to have a logistic growth pattern with initial population 1000 and an equilibrium population of 10,000. A count shows that at the end of 1 hr there are 2000 bacteria present. Determine the population as a function of time.
1
Expert's answer
2020-04-20T11:31:27-0400

"dN\/dt=aN(k-N)" is the logistic growth equation, where dN/dt is rate of growth of bacteria population, a is an undetermined constant and k is the saturation population or equilibrium population.

Solving the above equation we get;

"dN\/N(k-N)=adt=(1\/k)dN(1\/N+1\/(k-N))"

Integrating both sides we get;

"\implies \int dN(1/N+1/(k-N))=k\int adt"

"ln(N\/(k-N))=(ka)t+c"

Using the given conditions :

"k=10000; N(0)=1000; N(1)=2000"

We get;

"c=ln(1000\/(10000-1000))=-ln9=-2.197"

and

"ln(2000\/(10000-2000))=ka+c"

"\implies ka=(-ln4-c)=-1.386+2.197=0.811"


Thus, the population as a function of time is given by;

"ln(N\/(10000-N)))=0.811t-2.197"